# Fair Classifiers Without Fair Training: An Influence-Guided Data Sampling Approach

### v1

@@ Abstract
A fair classifier should ensure the benefit of people from different groups, while the group information is often sensitive and unsuitable for model training. Therefore, learning a fair classifier but excluding sensitive attributes in the training dataset is important. In this paper, we study learning fair classifiers without implementing fair training algorithms to avoid possible leakage of sensitive information. Our theoretical analyses validate the possibility of this approach, that traditional training on a dataset with an appropriate distribution shift can reduce both the upper bound for fairness disparity and model generalization error, indicating that fairness and accuracy can be improved simultaneously with simply traditional training. We then propose a tractable solution to progressively shift the original training data during training by sampling influential data, where the sensitive attribute of new data is not accessed in sampling or used in training. Extensive experiments on real-world data demonstrate the effectiveness of our proposed algorithm.
@@ 1 Introduction
Machine Learning (ML) has dramatically impacted numerous optimization and decision-making processes across various domains, including credit scoring (Siddiqi, 2005), demand forecasting (Carbonneau et al., 2008), and healthcare (Zicari et al., 2021). Algorithmic fairness embraces the principle, often enforced by law and regulations, that the decision-maker should not exhibit biases toward protected group membership (Zhu et al., 2022), identified by characteristics such as race, gender, or disability. The literature of fairness has gained prominence to mitigate the disparities in classification tasks. In essence, the core idea of in-processing mitigation algorithms is to impose some fairness constraints during the training process, leading to a widely embraced paradigm in practice (Donini et al., 2018; Hardt et al., 2016; Agarwal et al., 2018; Zafar et al., 2017; Wang et al., 2022a; Wei et al., 2023a). However, imposing fairness constraints can introduce two challenges in the realistic settings. First, the introduced parity constraint term relies on the visibility of sensitive attributes in the training dataset, which may be inaccessible due to privacy or safety concerns (Zhu et al., 2023). Second, enforcing fairness constraints or regularizers into machine learning models often leads to a tradeoff between fairness and accuracy (Kleinberg et al., 2016), underscoring a prevalent challenge: as we place greater emphasis on ensuring fairness, the predictive model performance might be weakened. With these two concerns in mind, our core inquiry rests upon the following question:
When sensitive attributes of training data are unknown and fair training algorithms are not employed, how can we develop a fair classifier without compromising the model accuracy?
This paper considers the following realistic scenario: we are given a large set of candidate samples without annotations. Due to the large cost of soliciting the true labels and privacy concerns of getting sensitive attributes, we are only allowed to access a small training dataset with labels and a small validation dataset with sensitive attributes. On the other hand, we are also curious about whether it is possible to learn the model without imposing any parity constraints, as constrained optimization will inevitably trade the accuracy rate and consume large computational resources. Concisely, our demand is a sampling algorithm that can effectively select a subset of training examples that suffices to mitigate the disparity without in-processing interventions.
Our proposal to address this problem is an influence-guided sampling strategy that actively samples new data points to supplement the training set. Our solution is motivated by a set of theoretical findings presented in Section 4. The analysis of the impact of data disparity on the group fairness metrics informs us that collecting new data to induce an appropriate distribution shift can effectively reduce fairness disparity without the need to implement any fair training algorithms on the updated training data.
The core challenge in crafting the sampling algorithm is approximating the influence of a candidate sample without having access to its sensitive attributes and true labels. Specifically, we assess the influence of each new unlabeled data example by comparing its gradient flow to that derived from the entire validation set, which quantifies the hypothesized change of parity metric when the specific example is included in the candidate training set. The requirement of using fairness constraints in training loss is relaxed by data sampling, as we found that the distribution shift incurred by appropriate data sampling may reduce the disparity without sacrificing model accuracy, even when the training loss does not contain any sensitive attributes. As a result, the sensitive attributes of the new unlabeled data are not required. Besides, before deciding to solicit the true label of an unlabeled example, we use the lowest-influencial label as a proxy for influence estimation. We name our solution Fair Influential Sampling (FIS).
The main contributions of our work are summarized as follows.
@@
    Our theoretical analysis indicates that training on datasets with a strategically implemented distribution shift can effectively reduce both the upper bound for fairness disparity and model generalization error (Lemma 4.1 and Theorem 4.2). This gives us a key insight that fairness and accuracy can be improved simultaneously even with simply traditional training. [Section 4]
@@
    We introduce a tractable solution (Algorithm 1) that progressively shifts the original training data, wherein we sample influential examples from an additional large dataset based on the combined influences of prediction and fairness without ever accessing their sensitive attributes during the training process. [Section 5]
@@ 2 Related Work
Active learning. The core idea of active learning is to rank unlabeled instances by developing specific significant measures, including uncertainty (Lewis and Catlett, 1994; Lindenbaum et al., 2004), representativeness (Dasgupta and Hsu, 2008), inconsistency (Wang et al., 2012), variance (Hoi et al., 2006), and error (Roy and McCallum, 2001). Each of these measures has its own criterion to determine the importance of instances for enhancing classifier performance. For example, uncertainty considers the most important unlabeled instance to be the nearest one to the current classification boundary.
A related line of work (Liu et al., 2021a; Wang et al., 2022b; Hammoudeh and Lowd, 2022; Xie et al., 2023) concentrates on ranking unlabeled instances based on the influence function. Compared to these studies with a focus on the prediction accuracy, our work poses a distinct challenge taking into account fairness violations. We note that adopting a particular sampling strategy can lead to distribution shifts between the training and testing data. What’s worse, even though fairness is satisfied within training dataset, the model may still exhibit unfair treatments on test dataset due to the distribution shift. Therefore, it becomes imperative for the sampling approach to also account for its potential impacts on fairness.
Fair classification. The fairness-aware learning algorithms, in general, can be categorized into pre-processing, in-processing, and post-processing methods. Pre-processing methods typically reweigh or distort the data examples to mitigate the identified biases (Asudeh et al., 2019; Azzalini et al., 2021; Tae and Whang, 2021; Sharma et al., 2020; Celis et al., 2020; Chawla et al., 2002; Zemel et al., 2013; Chen et al., 2018). More relevant to us is the importance reweighting, which assigns different weights to different training examples to ensure fairness across groups (Kamiran and Calders, 2012; Jiang and Nachum, 2020; Diesendruck et al., 2020; Choi et al., 2020; Qraitem et al., 2023; Li and Vasconcelos, 2019). Our proposed algorithm bears similarity to a specific case of importance reweighting, particularly the 0-1 reweighting applied to newly added data. Other parallel studies utilize importance weighting to learn a complex fair generative model in a weakly supervised setting (Diesendruck et al., 2020; Choi et al., 2020), or to mitigate representation bias in training datasets (Li and Vasconcelos, 2019).
Post-processing methods typically enforce fairness on a learned model through calibration (Feldman, 2015; Feldman et al., 2015; Hardt et al., 2016). Although this approach can decrease the disparity of the classifier, by decoupling the training from the fairness enforcement, this procedure may not lead to the best trade-off between fairness and accuracy (Woodworth et al., 2017; Pleiss et al., 2017). In contrast, these post-processing techniques necessitate access to the sensitive attribute during the inference phase, which is often not available in many real-world scenarios.
@@
Fairness-accuracy tradeoff. It has been demonstrated that there is an implicit trade-off between fairness and accuracy in the literature. Compared to the prior works (Menon and Williamson, 2018; Prost et al., 2019), our work does not require additional assumptions about the classifier and the characteristics of the training/testing datasets (for example, distribution shifts). Besides, the main difference is that this paper does not actually work in the classical regime of the fairness-accuracy tradeoff. By properly collecting new data from an unlabeled dataset, we can improve both generalization and fairness at the same time, which cannot be achieved by working on a fixed and static training dataset that naturally incurs such a tradeoff.
Li and Liu (2022) is a similar work that utilizes the influence function to reweight the data examples but requires re-training. Our method focuses on soliciting additional samples from an external unlabeled dataset while (Li and Liu, 2022) reweights the existing and fixed training dataset. Our approach is more closely with a fair active learning approach (Anahideh et al., 2022). However, this fair active learning framework relies on sensitive attribute information while our algorithm does not.
Distribution shifts. Common research concerning distribution shifts necessitates extra assumptions to build theoretical connections between features and attributes, like causal graphs (Singh et al., 2021), correlation shifts (Roh et al., 2023), and demographic shifts (Giguere et al., 2022). In contrast, our approach refrains from making further assumptions about the properties of distribution shifts.
In this literature, many works have utilized distributionally robust optimization (DRO) to reduce fairness disparity without sensitive attribute information (Hashimoto et al., 2018; Kirichenko et al., 2022; Liu et al., 2021b; Lahoti et al., 2020; Veldanda et al., 2023; Sohoni et al., 2020; Wei et al., 2023b). Although these works also evaluate the worst-group performance in the context of fairness, their approach differs as they do not strive to equalize the loss across all groups. Besides, in these studies, accuracy and worst-case accuracy are used to showcase the efficacy of DRO. Essentially, they equate fairness with uniform accuracy across groups, implying a model is considered fair if it demonstrates equal accuracies for all groups. However, this specific definition of fairness is somewhat restrictive and does not align with more conventional fairness definitions like DP or EOd.
3 Preliminary
3.1 Basic Scenario
We consider a standard $K$-class classification task, involving two distinct distributions: $\mathcal{P}$ for training (source) and $\mathcal{Q}$ for testing (target). Each distribution is defined as a probability distribution over certain examples, where each example, denoted as $z:=(x,y,a)$, comprises three random variables: the feature vector $x$, the label $y$, and the group membership $a$. The sensitive group $a$ often refers to characteristics such as race, gender, etc. Let $P$ and $Q$ represent the training and testing sets sampled from $\mathcal{P}$ and $\mathcal{Q}$, respectively. The training set $P$ is defined as $P:=\{(x_{n},y_{n},a_{n})\}_{n=1}^{N_{P}}$. Similarly, the test set is referred to as $Q:=\{(x_{n}^{\circ},y_{n}^{\circ},a_{n}^{\circ})\}_{n=1}^{N_{Q}}$. Note that, in practice, we are not privy to any details regarding the test dataset. Moreover, there might be potential distribution shifts between the train and test datasets, meaning that they are not independent and identically distributed (IID).
We are given a large unlabeled dataset $U:=\{(x^{\prime}_{n},\cdot,\cdot)\}_{n=1}^{N_{U}}$, which can be used for sociliting new annotations. In practice, the protected groups of datasets $P$ and $U$ remain undisclosed during training due to privacy and security concerns. The ultimate goal is to train a classifier $f$ trained on both $P$ and a small set of $U$ with a limited annotation budget for soliciting annotations while ensuring fairness constraints are satisfied on the validation dataset $Q_{v}$.
More formally, the learning objective can be formulated as
[…]
where $\mathbf{w}$ denotes the model parameters, $N_{r}(\ll N_{U})$ represents the annotation budget, and $\ell(\mathbf{w},z)$ represents the loss function on data point $z$, which can be an indicate function $\mathbb{I}(\cdot)$ to demonstrate 0-1 loss. $\phi(\mathbf{w},Q_{v})$ indicates the fairness violation on the validation dataset $Q_{v}$ with a tolerance threshold $\epsilon$. Here, we only require the validation data with the purpose of verifying the fairness constraints are satisfied, which is IID with the test dataset $Q$.
3.2 Fairness Definition
We then introduce the concerned fairness notions:
Definition 3.1.
@@ 3.2 Fairness Definition
(Risk disparity (Hashimoto et al., 2018; Zafar et al., 2019)). Define $Q_{k}$ as the subset of $Q$ corresponding to group $k$. Given the optimized model parameters $\mathbf{w}^{P}$ trained on set $P$, the risk disparity is defined as: $\mathcal{R}_{\mathcal{Q}_{k}}(\mathbf{w}^{P})-\mathcal{R}_{\mathcal{Q}}(\mathbf{w}^{P})$, where $\mathcal{R}_{\mathcal{Q}}(\mathbf{w}):=\mathbb{E}_{z\sim\mathcal{Q}}[\ell(\mathbf{w},z)]$ denotes the expected risk induced on target distribution $\mathcal{Q}$.
Definition 3.1 highlighted in (Hashimoto et al., 2018; Zafar et al., 2019) naturally quantifies the discrepancy in a trained model’s performance between a specific group set $Q_{k}$ and the entire test set $Q$. That is, a model can be deemed fair if it exhibits consistent performance for a group set $Q_{k}$ as compared to the test dataset $Q$. Even though there may be a general incompatibility between risk disparity and popular group fairness metrics like Demographic Parity (DP) and Equalized Odds (EOd), under the criteria of this definition, these metrics could be encouraged (Shui et al., 2022; Hashimoto et al., 2018). For example, in this context, demography parity can be defined as $\sum_{k,k^{\prime}}|\mathcal{R}_{\mathcal{Q}_{k}}(\mathbf{w}^{P})-\mathcal{R}_{\mathcal{Q}_{k}^{\prime}}(\mathbf{w}^{P})|$, measuring the summation of group disparities.
Remark 3.1.
@@
(Connections to other fairness definitions.) Definition 3.1 targets group-level fairness, which has similar granularity to the classical fairness notions such as the accuracy parity (Zafar et al., 2017), the device-level parity (Li et al., 2019), and small prediction loss for groups (Zafar et al., 2019; Balashankar et al., 2019; Martinez et al., 2019; Hashimoto et al., 2018).
4 Decoupling Fairness with Data Disparity
Including new data points to train fair classifiers raises a critical question: what strategy should we adopt to collect the data examples from the unlabeled dataset $U$ to mitigate fairness disparity without sacrificing accuracy? This task becomes particularly intricate without any further information about the training and testing data distributions, especially in the absence of sensitive attributes. In this section, we conduct a theoretical evaluation of the bounds of the risk disparity to ascertain the feasibility of our goal.
Without loss of generality, we discrete the whole distribution space and suppose that the train/test distributions are both drawn from a series of component distributions $\{\pi_{1},\cdots,\pi_{I}\}$ (Feldman, 2020). Then, the empirical risk $\mathcal{R}_{P}(\mathbf{w})$ calculated over a training set $P$ can be reformulated by splitting samples based on the component distributions:
$\mathcal{R}_{P}(\mathbf{w}):=$ $\mathbb{E}_{z\in P}[\ell(\mathbf{w},z)]=$ $\sum_{i=1}^{I}p^{(P)}(\pi=i)\cdot\mathbb{E}_{z\sim\pi_{i}}[\ell(\mathbf{w},z)].$ (2)
where $p^{(P)}(\pi=i)$ represents the frequencies of examples in $P$ drawn from component distribution $\pi_{i}$. Then, we can define the measure of probability distance between two sets or distributions as $\text{dist}(\mathcal{A},\mathcal{B}):=\sum_{i=1}^{I}|p^{(\mathcal{A})}(\pi=i)-p^{(\mathcal{B})}(\pi=i)|$.
To reflect the implicit unfairness in the models, we introduce two basic assumptions in convergence analysis (Li et al., 2019).
Assumption 4.1.
($L$-Lipschitz Continuous). There exists a constant $L>0$, for any $\mathbf{v},\mathbf{w}\in\mathbb{R}^{d}$, $\mathcal{R}_{P}(\mathbf{v})\leq\mathcal{R}_{P}(\mathbf{w})+\langle\nabla\mathcal{R}_{P}(\mathbf{w}),\mathbf{v}-\mathbf{w}\rangle+\frac{L}{2}\lVert\mathbf{v}-\mathbf{w}\rVert^{2}_{2}$.
Assumption 4.2.
(Bounded Gradient on Random Sample). The stochastic gradients on any sample $z$ are uniformly bounded, i.e., $\mathbb{E}[||\nabla\mathcal{R}_{P}(\mathbf{w}_{t},z)||^{2}]\leq G^{2}$, and training epoch $t\in[1,\cdots,T]$.
Analogous to Assumption 4.2, we further make a mild assumption to bound the loss over the component distributions $\pi_{i}$ according to the corresponding model, that is, $\mathbb{E}_{z\sim\pi_{i}}[\ell(\mathbf{w}^{P},z)]\leq G_{P},\forall i\in I$. For completeness, we first analyze the upper bound of generalization error, specifically from the standpoint of distribution shifts.
Lemma 4.1.
[…]
Note that the generalization error bound is predominantly influenced by the shift in distribution when we think of an overfitting model, i.e., the empirical risk $\mathcal{R}_{P}(\mathbf{w}^{P})\to 0$.
Theorem 4.2.
[…]
where $\Phi=4L^{2}G^{2}\sum_{t=0}^{T-1}(\eta_{t}^{2}(1+2\eta_{t}^{2}L^{2}))^{t},\Upsilon=\sqrt{\frac{\log(4/\delta)}{2N_{P}}}+\sqrt{\frac{\log(4/\delta)}{2N_{P_{k}}}}+\varpi+\varpi_{k}$.
Note that $\mathbb{E}_{z\sim\pi_{i}}[\ell(\mathbf{w}^{k},z)]\leq G_{k}$, $\varpi=\mathcal{R}_{P}(\mathbf{w}^{P})-\mathcal{R}_{\mathcal{Q}}^{*}(\mathbf{w}^{\mathcal{Q}})$ and $\varpi_{k}=\mathcal{R}_{P_{k}}(\mathbf{w}^{k})-\mathcal{R}_{\mathcal{Q}_{k}}^{*}(\mathbf{w}^{\mathcal{Q}_{k}})$. Specifically, $\varpi$ and $\varpi_{k}$ can be regarded as constants because $\mathcal{R}_{P}(\mathbf{w}^{P})$ and $\mathcal{R}_{P_{k}}(\mathbf{w}^{k})$ correspond to the empirical risks, $\mathcal{R}_{\mathcal{Q}}^{*}(\mathbf{w^{\mathcal{Q}}})$ and $\mathcal{R}_{\mathcal{Q}_{k}}^{*}(\mathbf{w^{\mathcal{Q}_{k}}})$ represent the ideal minimal empirical risk of model $\mathbf{w}^{\mathcal{Q}}$ trained on distribution $\mathcal{Q}$ and $\mathcal{Q}_{k}$, respectively.
Interpretations. The upper bound in Eq. (4) illustrates several intuitive aspects that induce unfairness. (1) Group biased data. For group-level fairness, the more balanced the data is, the smaller the risk disparity would be; (2) distribution shift. For source/target distribution, the closer the distribution is, the smaller the gap would be; (3) Data size. For training data size, the larger the size is (potentially eliminating data bias across groups), the smaller the gap would be.
Main observation. Note that Theorem 4.2 implies that the risk disparity is essentially influenced by both potential distribution shifts between the train and test datasets, as well as inherent group disparities. Our results provide guidance to reduce risk disparity by controlling the impact of distribution shifts via generalization error, which corresponds to practical model accuracy. That is, an appropriate distribution shift can reduce both the upper bound for risk disparity and generalization error. In an ideal scenario, the group gap in $P$ is smaller than the previous one in $Q$, and $\text{dist}(\mathcal{P},\mathcal{Q})$ is also small. Therefore, in general, distribution shifts typically with negative impacts on fairness can effectively be utilized to reduce the disparity without sacrificing accuracy, validating the possibility of our goal. More specifically, collecting new data pointrs to induce an appropriate distribution shift can effectively reduce disparities without the need to implement any fair training algorithms on the updated training data.
5 Improving Fairness with Data Influential Sampling
As motivated, in this section, we propose an influence-guided sampling strategy that actively solicits new data to supplement the train set, thus mitigating fairness disparity.

### v3

# Fairness Without Harm: An Influence-Guided Active Sampling Approach

@@ Abstract
The pursuit of fairness in machine learning (ML), ensuring that the models do not exhibit biases toward protected demographic groups, typically results in a compromise scenario. This compromise can be explained by a Pareto frontier where given certain resources (e.g., data), reducing the fairness violations often comes at the cost of lowering the model accuracy. In this work, we aim to train models that mitigate group fairness disparity without causing harm to model accuracy. Intuitively, acquiring more data is a natural and promising approach to achieve this goal by reaching a better Pareto frontier of the fairness-accuracy tradeoff. The current data acquisition methods, such as fair active learning approaches, typically require annotating sensitive attributes. However, these sensitive attribute annotations should be protected due to privacy and safety concerns. In this paper, we propose a tractable active data sampling algorithm that does not rely on training group annotations, instead only requiring group annotations on a small validation set. Specifically, the algorithm first scores each new example by its influence on fairness and accuracy evaluated on the validation dataset, and then selects a certain number of examples for training. We theoretically analyze how acquiring more data can improve fairness without causing harm, and validate the possibility of our sampling approach in the context of risk disparity. We also provide the upper bound of generalization error and risk disparity as well as the corresponding connections. Extensive experiments on real-world data demonstrate the effectiveness of our proposed algorithm. Our code is available at github.com/UCSC-REAL/FairnessWithoutHarm.
@@ 1 Introduction
Machine Learning (ML) has dramatically impacted numerous optimization and decision-making processes across various domains, such as credit scoring [62] and demand forecasting [14]. Algorithmic fairness embraces the principle, often enforced by law and regulations, that the decision-maker should not exhibit biases toward protected group membership [84], identified by characteristics such as race, gender, or disability. However, the pursuit of fairness unavoidably results in a compromise scenario where reducing the fairness violations usually leads to a degradation in accuracy, which has been observed and verified by numerous literature [52, 25, 81, 69, 83]. Theoretically, the phenomenon can be understood through a Pareto frontier on the tradeoff between group fairness and accuracy [72, 51, 10, 74]. That is, as illustrated in Figure 1, given certain resources such as training data, when a model has reached a point on the Pareto frontier, without more data resources, it is impossible that one can improve fairness without worsening off model accuracy.
One major source of unfairness and a major cause of the fairness-accuracy tradeoff is biased training data. If an unbiased and “fairer” dataset is available, we will be hopeful that unfairness can be alleviated without compromising accuracy. Furthermore, such a “fairer” dataset would allow for obtaining a fair and accurate model through the standard empirical risk minimization (ERM) with cross-entropy (CE) loss. The above observation points to a promising way to improve fairness via actively acquiring more informative data, aiming to shift towards a better Pareto frontier of the fairness-accuracy trade-off [1, 46]. However, existing approaches that seek more data, such as fair active learning [4], typically require annotating sensitive attributes for training data. In practice, these sensitive attribute information such as race and gender, should be protected due to privacy regulations [35, 5, 66]. In the normal active learning scenario, collecting more data with sensitive attributes heightens privacy and safety risks due to the increased probability of leaking sensitive information.
Figure 1: We compare the Pareto frontiers between the model trained with scarce data and that trained with rich data. Acquiring more data is capable of shifting the Pareto frontier toward lower disparity and lower error rates. In consequence, we can reach a new trade-off point that offers improved fairness and accuracy simultaneously, surpassing the original trade-off point.
Therefore, we ask the following question: When not disclosing more annotations of training sensitive attributes, how can we acquire more data to improve model fairness without sacrificing accuracy?
In this paper, we propose a tractable active data sampling algorithm in a training sensitive attributes-free way, which solely requires sensitive attributes on a small validation set. Particularly, the algorithm evaluates each example’s influence on fairness and accuracy using the validation dataset for ranking and then selects a certain number of examples to supplement the training set for training. We name our solution Fair Influential Sampling (FIS). The core challenge is approximating the corresponding influences of each new example without accessing its sensitive attributes. Technically, we evaluate the importance (influences) of each new example by comparing its gradient to that derived from the entire validation set. This comparison helps quantify the hypothesized change of group fairness disparity metric when adding this example to the training set. As a result, the requirement of training sensitive attributes can be relaxed, as gradient derivation serves as a role of fairness constraints to measure the group fairness disparity. The main contributions of our work are summarized as follows.
@@
    We develop a tractable active data sampling algorithm (Algorithm 1) that does not rely on training sensitive attributes. The algorithm scores each new example based on the combined influences of prediction and fairness and then opts for a certain number of examples for training. [Section 4]
@@
    We theoretically analyze how acquiring more data can improve fairness without harm from a distribution shift perspective view, and validate the possibility of our sampling approach in the context of risk disparity. We also provide the upper bound of generalization error and risk disparity as well as the corresponding connections (Theorem 5.1 and Theorem 5.2). [Section 5]
@@ 2 Related work
Fairness-accuracy tradeoff
@@
There are numerous works that have been successful at mitigating fairness disparities [24, 32, 2, 78, 69]. However, these works typically rely on protected sensitive attributes of training examples to measure the fairness disparities across groups. Moreover, a fairness-accuracy tradeoff has been shown, meaning that enforcing fair constraints heavily degrades the model performance [52, 25, 81]. Notably, Chen et al. [20] characterized the change of the fairness violation when the data distribution is shifted. Except for training sensitive attributes, this paper does not work in the classical regime of the fairness-accuracy tradeoff. By properly collecting new data, we can improve both accuracy and fairness, which cannot be achieved by working on a static training dataset that naturally incurs such a tradeoff. Besides, compared to prior works [52, 56], our method does not require additional assumptions about the classifier and the characteristics of the training/testing datasets (e.g., distribution shifts). Relevant work [42] utilizes the influence function to reweight the data examples but requires re-training. Our method focuses on soliciting additional samples from an external dataset while [42] reweights the existing and fixed training dataset.
Active learning
The core idea of active learning is to rank unlabeled instances by developing specific measures, including uncertainty [41, 45], representativeness [22], inconsistency [70], variance [34], and error [59]. A related line of work [48, 71, 31, 76] concentrates on ranking unlabeled instances based on the influence function. Compared to these studies with a focus on prediction performance, our work poses a distinct challenge taking into account fairness violations. Our approach is more closely with the fair active learning approach [4]. However, this framework still relies on training sensitive attributes and then unavoidably encounters the tradeoff between fairness and accuracy.
Fair classifiers without demographics
There are various studies to achieve fairness without demographics. For example, Zhao et al. [82] explores the correlations between sensitive attributes and non-sensitive attributes to learn fair and accurate classifiers. Yan et al. [77] investigates the class imbalance problem with a KNN-based pre-processing method. Chai et al. [16] utilizes the soft labels from an overfitting teacher model to train a student model to avoid using demographics. A line of research establishes theoretical connections between features and attributes to avoid using demographic information, employing methods like causal graphs [63], correlation shifts [58], and demographic shifts [30]. In contrast, our approach refrains from making assumptions. Another line of work utilizes distributionally robust optimization (DRO) to reduce fairness disparity without relying on training sensitive attributes [33, 38, 47, 40, 67, 64]. Although these works evaluate the worst-case group performance in the context of fairness, their approaches differ as they do not strive to equalize the loss across groups. Besides, in these studies, accuracy and worst-case accuracy are used as fairness metrics to showcase the efficacy of the proposed algorithms. However, these fairness metrics are restrictive and inconsistent with common definitions such as demographic parity (DP).
Fair classification
The fairness-aware learning algorithms, in general, can be categorized into pre-processing, in-processing, and post-processing methods. Pre-processing methods typically reweigh or distort the data examples to mitigate the identified biases [7, 9, 65, 60, 15, 17, 80, 18, 68]. More relevant to this work is the importance reweighting, which assigns weights to training examples [37, 36, 23, 21, 57, 44]. Our algorithm bears similarity to a specific case of importance reweighting, particularly the 0-1 reweighting applied to newly added data. Other parallel studies utilize importance weighting to learn a complex generative model in a weakly supervised setting [23, 21], or to mitigate representation bias in training datasets [44]. Post-processing methods typically enforce fairness on a learned model through calibration [26, 27, 32], but these work might not achieve the best fairness-accuracy tradeoff [75, 55]. In contrast, these post-processing works still require sensitive attributes during the inference phase. Recent work [19] develops a bias score classifier that operates independently of sensitive attributes; however, it is constrained to binary classifications.
3 Preliminaries
Problem setup
We consider a standard $K$-class classification task whose training (test) data distribution is $\mathcal{P}$ ($\mathcal{Q}$). Let $P:=\{z_{n}\}_{n=1}^{|P|}$ represent the training dataset following distribution $\mathcal{P}$, where $|P|$ denotes the corresponding sample size. Each example, denoted as $z_{n}:=(x_{n},y_{n})$, comprises two random variables: the feature vector $x$ and the label $y$. The model trained on $P$ is evaluated by the test dataset $Q:=\{z^{\circ}\}_{n=1}^{|Q|}$, where ${(\cdot)}^{\circ}$ denotes that the data follows distribution $\mathcal{Q}$, each example $z_{n}^{\circ}:=(x^{\circ},y^{\circ},s^{\circ})$, and the sensitive group $s_{n}^{\circ}$ often refers to characteristics such as race, gender, etc. To align the fairness requirements on the test set $Q$ with the model trained on $P$, a popular way is to exploit the sensitive attributes $s$ [24, 72] or their proxies [84] in $P$ and use them to formulate Lagrangians during training. However, extending these approaches to the active learning setting would require disclosing more sensitive attributes during sampling and training [4], which contradicts our goal. To avoid disclosing more sensitive attributes, we align the fairness requirements by a small hold-out validation dataset $Q_{v}:=\{z_{n}^{\circ}\}_{n=1}^{|Q_{v}|}$ that is independent and identically distributed (IID) with the test set $Q$. We defer more technical details to Section 4.
Following the general active learning setting [4], we would acquire new examples from a large unlabeled dataset $U:=\{x^{\prime}_{n}\}_{n=1}^{|U|}$ within a limited labeling budget $B(\ll|U|)$ [48, 71, 31]. Denote the solicited example by $z^{\prime}_{n}:=(x^{\prime}_{n},y^{\prime}_{n})$, where $y^{\prime}_{n}$ is the ground-truth label. Note that the protected sensitive attributes from datasets $P$ and $U$ remain undisclosed during sampling and training. In this paper, we aim to incrementally update a model that was initially trained on $P$ using standard ERM, by incorporating newly solicited data $z_{n}^{\prime}$, such that the model can improve fairness without worsening model accuracy. Thus, the core challenge is efficiently determining new examples that induce a significantly better Pareto frontier. In the proceeding section 4, we shall delve into how to acquire new data from unlabeled set.
Fairness definition
Note that this work focuses solely on active sampling to build a fairer dataset, which is then used to train the model through standard ERM with Cross-Entropy (CE) loss. Without relying on additional assumptions about the model or training/testing dataset, an intuitive and natural approach is to analyze the expected risk. Therefore, we introduce the concept of risk disparity as an intermediate-term for theoretical analysis of fairness.
Definition 3.1 (Risk disparity [33, 79, 3]).
@@ 3 Preliminaries
Define $Q_{k}$ as the sub-distribution of $Q$ corresponding to group $k$. Given the optimized model parameters $\mathbf{w}^{P}$ trained on set $P$, risk disparity is defined as: $\mathcal{R}_{\mathcal{Q}_{k}}(\mathbf{w}^{P})-\mathcal{R}_{\mathcal{Q}}(\mathbf{w}^{P})$, where $\mathcal{R}_{\mathcal{Q}}(\mathbf{w}):=\mathbb{E}_{z\sim\mathcal{Q}}[\ell(\mathbf{w},z)]$ denotes expected risk induced on target distribution $\mathcal{Q}$.
Definition 3.1 naturally quantifies the discrepancy in a trained model’s performance between a specific group set $Q_{k}$ and the entire test set $Q$. That is, a model can be deemed fair if it exhibits consistent performance for a group set $Q_{k}$ as compared to the test dataset $Q$. In settings such as face or speech recognition, this fairness definition implies the necessity for all demographic groups to receive the same quality service [13]. For completeness, we also include two well-known definitions of fairness:
Definition 3.2 (Demographic Parity (DP)).
A classifier $f$ adheres to demographic parity concerning the sensitive attribute $s$ if: $\mathbb{E}[f(\mathbf{w},x)]=\mathbb{E}[f(\mathbf{w},x)|s]$.
Definition 3.3 (Equalized Odds (EOd [32]).
A classifier $f$ meets the equalized odds with respect to the sensitive attribute $s$ if: $\mathbb{E}[f(\mathbf{w},x)|y]=\mathbb{E}[f(\mathbf{w},x)|y,s]$.
Even though there may be a general incompatibility between risk disparity and popular group fairness metrics like DP and EOd, under the criteria of the proposed fairness notion, these definitions could be encouraged [61, 33]. More details and proof can be found in the Appendix B.
Proposition 3.1.
(Informal) Under appropriate conditions, the risk disparity can serve as a lower bound for fairness disparities based on common fairness definitions, such as DP and EOd.
Remark 3.1 (Connections to other fairness definitions).
@@
Definition 3.1 targets group-level risk fairness, which has similar granularity to other fairness notions such as accuracy parity [78], device-level parity [43], small accuracy loss for groups [79, 10, 50, 33], and bounded group loss [3].
4 Improving fairness without harm via data influential sampling
In this section, we first introduce how to measure the importance (influence) of each example on accuracy and fairness without using the corresponding sensitive attributes, respectively. Then, we propose an influence-guided sampling algorithm that actively acquires new data based on the influences for further training.